OpenStudy (anonymous):

Please explain this to me!! An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

@agent0smith @Studygirl14 @mathmale @ranga @patrickStar @MayankD

OpenStudy (anonymous):

still need help??

OpenStudy (anonymous):

yes severly!!

OpenStudy (anonymous):

isolate y, what do you have??

OpenStudy (anonymous):

-y=-2-9x??

OpenStudy (anonymous):

take off all minus signs, you have y = 9x+2, ok??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so do i plug in y to the equation??

OpenStudy (anonymous):

So, for part A) solution set ={(x,y) | y = 9x +2, x, y in R} That is NOTATION. "|" means "such that" Verbalize it, it will be said: "solutions are set of point (x,y) such that y =9x+2, x and y are in real "

OpenStudy (anonymous):

Part B), just arbitrary put x to get y you can put x =0, then y = 2, so you have (0,2) is one point you can put x =1, then y =..... to have another point you can put x = whatever , then y = .... to have one another point and ..... ANSWER hihihihi

OpenStudy (anonymous):

Got me so far??

OpenStudy (anonymous):

trying to just give me a minute to process it:)

OpenStudy (anonymous):

ok, take your time. :)

OpenStudy (anonymous):

ok so i could put x = 1 and y=3 and x=2 and y=4 for part B... Right??

OpenStudy (anonymous):

(0,2) (1,3) (2,4)

OpenStudy (anonymous):

hey, x =1, y\(\neq 3\), let plug y = 9*x +2 , plug x =1into y , then, y =9*1+2 =11\(\neq 3\)

OpenStudy (anonymous):

\[y=9*\color{red}{x}+2\\y=9*\color{red}{1}+2=11\] got what I meant?

OpenStudy (anonymous):

yah i think!!! y=9(2)+2 -> 18+2 --> 20....so it would be (2,20) for my last one??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that was for part B right?? the 3 different solutions??

OpenStudy (anonymous):

yeah, but I don't understand what part C is. @agent0smith Please

OpenStudy (anonymous):

A system of equations is "inconsistent" if no solutions exist. Graphically, this means that two lines are not intersecting. Can you draw a line that does not intersect that graph of our above equatioN????

OpenStudy (anonymous):

i read this but doent truly understand it...

OpenStudy (anonymous):

if there are 3 solutions on the graph they are bound to intersect right?? (0,1) (1,11) (2,20) will intersect right??

OpenStudy (anonymous):

2 lines are not intersecting iff they are parallel!!

OpenStudy (anonymous):

in R^2, that is only 1 case to get 2 lines are not intersecting (that 's they are parallel) if it is the case, then just replace the last number I mean: y = 9x + whatever ( not 2) you get the answer you can pick any number \(\neq 2\) which will make the new one is parallel to the given one.

OpenStudy (anonymous):

so... y=9x+5 would work and just put that as the equation that can be paired to form a system of equations that is inconsistent.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Oh my gosh thanks so much!!! how do you give a medal this is my first time on this thing.

OpenStudy (anonymous):

NOTATION again, hihihi \[\begin{cases} y=9x+2\\y =9x+5\end{cases}\]

OpenStudy (anonymous):

oh this is hannah not nate i dont want my brother sounding like a girl:)

OpenStudy (anonymous):

no need to give me medal. I like students who is willing to study, not just ask for the answer. hihihi

OpenStudy (anonymous):

ok:) i dont even know what a medal would be for but ive seen others do it:) thanks again!!!